# Non-Vacuous Generalization Bounds for Large Language Models

### 28 Dec 2023 v1

@@ 4.3 Accommodating the Unbounded NLL Objective Using Prediction Smoothing
where $p_{\theta}(x_{i}|x_{<i})$ is the base model of token probabilities, $\alpha\in(0,1)$ is the mixing parameter, and $p_{h}(x_{i}|x_{<i})$ is the smoothed predictor. The model on an entire sequence $X$ is defined autoregressively in terms of this mixture model $p_{h}(X):=\Pi_{i}^{L}p_{h}(x_{i}|x_{<i})$, and we find this to be a more effective way of constructing the bounds than constructing the mixture at the sequence level. In analogy to label smoothing where the labels of the training objective are mixed with the uniform distribution, we term this operation as prediction smoothing.
Notably, while the predictions $p_{h}(x_{i}|x_{<i})$ are not independent, the predictions on an entire sequence $p_{h}(X)$ are independent for sequences $X$ which occupy different non-overlapping context windows of the model. Therefore with the BPD evaluated on a single context chunk of length $L$, $\mathrm{BPD}(h,X):=-\log_{2}p_{h}(X)/L$, we define the empirical risk as $\hat{R}(h)=\frac{1}{m}\sum_{k=1}^{m}\mathrm{BPD}(h,X_{k})$, the average over independent chunks $\{X_{k}\}_{k=1}^{m}$. With this construction, the BPD for sequence $X$ can be bounded as follows:
$\log_{2}(V/\alpha)-\log_{2}\big{(}1+(1-\alpha)V/\alpha\big{)}\leq\mathrm{BPD}(h,X)\leq\log_{2}(V/\alpha),$ (4)
as we show in Section A.2. Using $\Delta=\log_{2}\big{(}1+(1-\alpha)V/\alpha\big{)}$, we can plug these values into Equation 1 to generate bounds for the number of bits per dimension. We explore the trade-off over different values of $\alpha$ in Figure 2 (right). As $\alpha$ gets larger, the interval size $\Delta$ representing the worst case behavior goes down, whereas the empirical risk goes up, leading to a sweet spot in the middle. By defining the hypothesis $h=(\theta,d,r,\alpha)$ to include the model parameters, LoRA space hyperparameters $d,r$, and the mixture weight $\alpha$, we can view $\alpha$ as merely one additional model parameter accounted in $\log 1/P(h)$. By doing so, we are free to optimize over $\alpha$ in the computation of the bound, and we can do so without retraining the model.
To satisfy the i.i.d assumption in practice, we divide a large text corpus into non-overlapping chunks of size equal to a context length. The dataset is made up of these chunks, so that a single sample from the dataset includes all of the tokens in the given chunk. Then, we draw i.i.d samples from the uniform distribution over this dataset. It is important to note that even when the chunks in the dataset have relationships with each other, drawing i.i.d samples ensures that we satisfy the assumptions of finite hypothesis bounds. Unlike some models which draw on the previous history that lies outside the current context window, such as with Transformer-XL (Dai et al., 2019) or Mistral-7B (Jiang et al., 2023), the GPT-2 architecture (Radford et al., 2019) that we use in this work considers only the context and not any previous history.
Our bounds for non-overlapping sequences are still significant since the bits-per-dimension for a given sequence can be computed as the average error for each token in the sequence given previous tokens, where the token error here refers to the negative log-likelihood $\mathrm{BPD}(h,X):=-\log_{2}p_{h}(X)/L=-\sum_{i}^{L}\log_{2}p_{h}(x_{i}|x_{<i})/L$. Therefore, an upper bound on the expected BPD error still reflects a guarantee on the average performance of the model at the token level, conditioned on previous tokens within independent sequences, and is a common quantity of interest in language modeling.
@@ 4.4 Using Subsampling in Bound Computation
The empirical risk requires evaluating the model on the entire training dataset of $m$ data points: $\hat{R}(h)=\frac{1}{m}\sum_{i=1}\hat{R}_{i}(h)$. As large language models are typically trained for only $1$ epoch or less, doing so is prohibitively expensive. Instead, we propose to modify our generalization bounds to account for evaluating only a subsample of size $n\ll m$ of the training dataset when computing the empirical risk.
@@
where $s=n/(n+m)$.
Using this subsampling bound, we can get massive savings in the cost of computing a bound for a given model. For dataset sizes in the $10$’s of millions, we can get away with evaluating only $10$k data points after the model has been trained, with a negligible penalty in the bounds. In fact, we need not even train on the entirety of the training data in order to produce valid bounds. For LLMs where the cost of training is extremely high, we may wish to terminate the training of the compressed model much earlier than would be optimal for models that get deployed.
@@ 5 Non-Vacuous Generalization Bounds for LLMs
We outline the pretraining and bound computation pipeline and then present our empirical results.
End-to-end pipeline. Assembling the components described in Section 4, we train variants of a GPT-style architecture through the non-linear compressed parameterization in Equation 2. We use several values for the subspace dimension $d$ and two values for the rank of the LoRA matrices $r$. Nearing the end of training, we train for additional steps using quantization-aware training with a small number of quantization levels (with additional details listed in Appendix C). We express $w$ in this quantization and encode it using arithmetic coding to determine the compressed size of the model. Added to the size of the model are the bits needed to encode the choice of $d,r$, $\alpha$, the learning rate, and the quantization levels. We then evaluate the empirical log probabilities and token predictions for each token in the sequence on a small subset of the training data $n=10000$. With these predictions, we can compute the generalization bound in Equation 5 as a function of $\alpha$, and we optimize over this parameter for each model. Finally, we can tune the extent of compression through the different choices of $d$ and choose the subspace dimension that produces the best bound.
For the bound computation in this section, we consider a GPT-style architecture and use the OpenWebText dataset to pretrain it from scratch on next token prediction using SubLoRA. We use our pipeline to compute the generalization bounds and report the results in Table 1. We consider the token level error averaged over the sequence as the empirical risk which we bound. For instance, the Top-1 Error Bound refers to the upper bound on the expected Top-1 error per token averaged over the chunk $R(h,X_{k})=\frac{1}{L}\sum_{i=1}^{L}\mathbf{1}{[\operatorname*{arg\,max}p(x_{i}|x_{<i}=x_{<i}^{k})=x_{i}^{k}]}$, where the upper index $k$ denotes the chunk index and the lower index denotes the position within the chunk.
The best bound is indeed obtained by using our SubLoRA compression technique, which combines the strengths of both low rank adaptation and subspace training. When we solely apply quantization and arithmetic coding without implementing LoRA or linear subspace compression during the training phase, we obtain vacuous bounds.
Table 1: Our best generalization bounds achieved for the GPT-2 architecture for BPD and Top-k token prediction error, all of which are non-vacuous.
@@
Top-1 Error (%) $\mathbf{96.17}$ 100 97.40 100 99.99
Top-10 Error (%) $\mathbf{78.18}$ 85.85 80.15 100 99.98
Top-100 Error (%) $\mathbf{58.72}$ 65.19 76.11 100 99.80
Bits per Dimension $\mathbf{12.09}$ 12.90 14.68 65.37 15.62
@@ 6 Understanding the Generalization of LLMs
As language models grow in size, it is clear that they gain an increasing capacity to fit their training data. On the one hand, this increasing capacity might mean that, as LLMs become capable of learning increasingly complex functions, they become increasingly likely to merely memorize their training samples and not perform any meaningful generalization beyond their training corpora. After all, they have many more parameters to use in fitting the data. On the other hand, large language models have proven to be surprisingly capable at generalizing, often extending to tasks that seem quite different from the training objective.
We investigate the tension between these two narratives along several fronts: We assess how generalization bounds change with the size of the model, whether language models can form a compression of the training data even when accounting for their large size, and how structure in the training data affects the generalization of the learned model.
@@ 6.1 Larger Models Are More Compressible and Generalize Better
We evaluate our bounds on a collection of LLMs with different numbers of parameters, choosing the appropriate scaling for the width, depth, number of attention heads etc. Surprisingly, we find that our generalization bounds in fact improve with model size, even as the training dataset is held fixed. With our SubLoRA compression, larger models are more compressible given a fixed training error. These results are shown in Figure 3. While some explanations for why larger models should generalize better have been put forward in the literature (Nakkiran et al., 2021; Gunasekar et al., 2017), the mechanism by which larger models become more compressible is not clear, and we believe this result is noteworthy and requires further investigation.
We also note that in addition to constructing generalization bounds, we can use our compressed models to form a compression of the training dataset itself. In Figure 3, we count the number of bits needed to encode the model $C(h)$ and the number of bits to encode the data using the model $C(\{X\}_{i=1}^{m}|h)$, which is the negative log likelihood of the entire dataset according to the model. Adding these two up, we have a compression of the training dataset using the model, and one which is closely related to our generalization bounds.
@@ 6.2 How does Generalization of LLMs Depend on Structure in Text?
Neural networks that fit a training dataset of random noise will not be able to generalize, and the ability of overparametrized networks to fit noise implies that uniform convergence is impossible across the general hypothesis class (Nagarajan & Kolter, 2019). This fact is a clear demonstration that the structure of the dataset influences the generalization properties of the model. However, the impact of more subtle structures on generalization is less understood theoretically. Hence, we use our bounds to investigate how the temporal order structure relates to compressibility and generalization.
We train models that explicitly break the temporal structure of the text data by applying random permutations to each sequence during training. Consequently, the model can only make use of the input information as if it were a bag of words. We find that this broken order structure indeed leads to less favorable generalization bounds. Figure 4 shows the best error bounds when the original and perturbed data are used to train the model and evaluate the bounds for the bits per dimension, top-1 error and top-100 error losses. While the top-1 error bound becomes vacuous as we break the text structure, the top-100 error and bits per dimensions bounds remain non-vacuous. This might be due to the fact that as we perturb the sequence, predicting the next token accurately becomes an extremely difficult task for LLMs, while predicting a token that fit generally into the context, without necessarily being the correct token, is an easier task.
7 Discussion and Future Directions
Despite containing a very large number of parameters, we demonstrated that large language models are highly compressible. We further provided the first non-vacuous generalization bounds for LLM pretraining by using extreme levels of model compression. Our bounds suggest that compression bounds present new possibilities for understanding how and why language models generalize.
We close with a discussion of the limitations of this work, along with their implications for future generalization theory of language models:
@@ 7 Discussion and Future Directions
Appendix
@@ A.1 Finite Hypothesis Bound
Consider a bounded risk $R(h,x_{i})\in[a_{i},a_{i}+\Delta_{i}]$ and a finite hypothesis space $h\in\mathcal{H}$ for which we have a prior $P(h)$. Let the empirical risk $\hat{R}(h)=\frac{1}{m}\sum_{i=1}^{m}R(h,x_{i})$ be a sum over independent random variables $R(h,x_{i})$ for a fixed hypothesis $h$. Let $R(h)=\mathbb{E}[\hat{R}(h)]$ be the expected risk and $\hat{\Delta}=\sqrt{\frac{1}{m}\sum_{i=1}^{m}\Delta_{i}^{2}}$.
@@
$R(h)\leq\hat{R}(h)+\hat{\Delta}\sqrt{\frac{\log 1/P(h)+\log 1/\delta}{2m}},$ (A.1)
@@
$t(h)=\hat{\Delta}\sqrt{\frac{\log 1/P(h)+\log 1/\delta}{2m}}$ (A.2)
@@
$R(h)\leq\hat{R}(h)+\hat{\Delta}\sqrt{\frac{\log 1/P(h)+\log 1/\delta}{2m}}.$

### 13 Feb 2024 v2

@@ 4.3 Accommodating the Unbounded Negative Log-Likelihood
where $p_{\theta}(x_{i}|x_{<i})$ is the base model of token probabilities, $\alpha\in(0,1)$ is the mixing parameter, and $p_{h}(x_{i}|x_{<i})$ is the smoothed predictor.
The model on an entire document $X=\{x_{i}\}_{i=1}^{L}$ composed of $L$ tokens is defined autoregressively in terms of this mixture model $p_{h}(X):=\Pi_{i}^{L}p_{h}(x_{i}|x_{<i})$, and we find this to be a more effective way of constructing the bounds than constructing the mixture at the document level. In analogy to label smoothing where the labels of the training objective are mixed with the uniform distribution, we term this operation as prediction smoothing.
As we show in Section A.2, the NLL of the prediction smoothed model on a document $\mathrm{BPD}(h,X):=-\log_{2}p_{h}(X)/L$ can be bounded as follows:
$\log_{2}(V/\alpha)-\Delta$ $\leq\mathrm{BPD}(h,X)\leq\log_{2}(V/\alpha),$
for $\Delta=\log_{2}\big{(}1+(1-\alpha)V/\alpha\big{)}$. With prediction smoothing, the risk $R(h,X)=\mathrm{BPD}(h,X)$ on a given document is bounded in an interval of size $\Delta$, and therefore we can use Equation 1 to generate bounds for negative log-likelihood of this model. We refer to $\Delta$ as the worst-case interval size.
We explore the trade-off over different values of $\alpha$ in Figure 2 (right). As $\alpha$ gets larger, the interval size $\Delta$ representing the worst-case behavior goes down, whereas the empirical risk goes up, leading to a sweet spot in the middle. By defining the hypothesis $h=(\theta,d,r,\alpha)$ to include the model parameters, LoRA space hyperparameters $d,r$, and the mixture weight $\alpha$, we can view $\alpha$ as merely one additional model parameter accounted in $\log 1/P(h)$. By doing so, we are free to optimize over $\alpha$ in the computation of the bound, and we can do so without retraining the model.
@@ 4.4 Using Subsampling in Bound Computation
The empirical risk requires evaluating the model on the full training dataset of $m$ data points: $\hat{R}(h)=\frac{1}{m}\sum_{i=1}\hat{R}_{i}(h)$. As large language models are typically trained for only $1$ epoch or less, doing so is prohibitively expensive. Instead, we propose to modify our generalization bounds to account for evaluating only a subsample of size $n\ll m$ of the training dataset when computing the empirical risk.
@@
where $s=n/(n+m)$. Using this subsampling bound, we can accelerate bound computation. For dataset sizes in the $10$’s of millions, we can get away with evaluating only $10,000$ data points after the model has been trained, with a negligible penalty in the bounds. In fact, we need not even train on the entirety of the training data in order to produce valid bounds as long we indeed sample uniformly.
5 SubLoRA: A Simple and Efficient Nonlinear
Parameterization of the Hypothesis Space
To find compressible solutions $h$ that simultaneously are expressive enough to achieve low training error, we search over a carefully designed manifold of possible parameters that live within the parameter space.
In contrast to Lotfi et al. (2022), we consider a nonlinear parameterization of the model weights $\theta=f(\theta_{0},w)$ given by the composition of LoRA (Hu et al., 2021) (a nonlinear parameterization) and the subspace compression matrices. Given a vector of model parameters $\theta$, we break down its constituent components into the different weight matrices $W_{i}$ and associated biases $b_{i}$: $\mathrm{unflatten}(\theta)=\{(W_{i},b_{i})\}_{i\in I}$. We define a nonlinear parameterization of the hypothesis space,
$\theta=\theta_{0}+\mathrm{LoRA}(Pw),$ (4)
where $\mathrm{LoRA}$ is defined by the implementation of the low-rank products for the weight matrices, leaving the biases unchanged. As $Pw$ and $\theta$ are the flattened parameter vectors, $\mathrm{LoRA}(\cdot)$ is defined as the operation that unflattens the vector, applies the low-rank product, and then flattens the result. Here, $\theta_{0}$ is merely a random initialization of the model parameters, and $P\in\mathbb{R}^{D\times d}$ is a Kronecker product projector $P=Q_{1}\otimes Q_{2}$ for $Q_{1},Q_{2}$ constructed by orthogonalizing Gaussian random matrices by QR factorization: $P_{1},P_{2}\sim\mathcal{N}(0,1/\sqrt{D})^{\sqrt{D}\times\sqrt{d}}$ with $Q_{1}R_{1}=P_{1}$ and similarly for $Q_{2}$. We apply LoRA only over the self-attention layer and the last linear layer weight matrices, meaning that other model parameters do not differ from their initialized values. While LoRA was developed for finetuning LLMs, we find that even when pretraining using LoRA, we can achieve non-trivial performance. In order to compress the model, we need only to represent the vector $w$ since $\theta_{0}$ and $P$ are chosen ahead of time and specified in the architecture via random initialization.
In Figure 1 (left), we show the Pareto frontier of empirical risk and the complexity penalty in the relevant generalization bound with LoRA, subspace training, and SubLoRA. Rather than being competing methods for compression, LoRA and subspace training are complementary and exploit different structures in the parameter space to provide a family of models in the original hypothesis space that are both expressive and compressible. SubLoRA achieves a strict improvement over LoRA and subspace training, often being the deciding factor whether the bounds are vacuous or non-vacuous. In Figure 2 (left), we explore how the compressed size of the model and the empirical risk vary as a function of the subspace dimension $d$.
@@ 6 Non-Vacuous Generalization Bounds for LLMs
We outline the pretraining and bound computation pipeline and present our empirical results.
6.1 End-to-end Pipeline
Assembling the components described in Section 4, we train variants of a GPT-style architecture through the nonlinear compressed parameterization in Equation 4. We use several values for the subspace dimension $d$ and two values for the rank of the LoRA matrices $r$. Nearing the end of training, we train for additional steps using quantization-aware training with a small number of quantization levels (with additional details listed in Appendix D). We express $w$ in this quantization and encode it using arithmetic coding to determine the compressed size of the model. Added to the size of the model are the bits needed to encode the choice of $d,r$, $\alpha$, the learning rate, and the quantization levels.
We evaluate the empirical log probabilities and token predictions for each token in the sequence on a small subset of the training data $n=10000$. With these predictions, we can compute the generalization bound in Equation 3 as a function of $\alpha$, and we optimize over this parameter for each model. Finally, we can tune the extent of compression through the different choices of $d$ and choose the subspace dimension that produces the best bound.
6.2 Non-Vacuous Bounds for GPT-2 Small
We consider the GPT-2 small architecture with $124$M parameters and compute our next token prediction document-level bounds by pretraining these models on the OpenWebText dataset using SubLoRA. We report the results in Table 1. We consider the token level error averaged over a document as the empirical risk. For instance, the Top-1 Error Bound refers to the upper bound on the expected Top-1 error per token averaged over the document $R(h,X_{k})=\frac{1}{L}\sum_{i=1}^{L}\mathbf{1}{[\operatorname{argmax}p(x_{i}|x_{<i}=x_{<i}^{k})=x_{i}^{k}]}$, where the upper index $k$ denotes the document index and the lower index denotes the position within the document. Random guess performance is $\log_{2}V$ for BPD and $1-k/V$ for Top-k Error.
The best bounds are indeed obtained using our simple compression technique, which combines the strengths of both low-rank adaptation and subspace training. When we solely apply quantization and arithmetic coding without implementing LoRA or linear subspace compression during the training phase, we obtain vacuous bounds.
Note 0 (Significance of our bounds with I.I.D sampling).
The bits-per-dimension for a given document can be computed as the average error for each token in the sequence given previous tokens withing the same document, where the token error here refers to the negative log-likelihood $\mathrm{BPD}(h,X):=-\log_{2}p_{h}(X)/L=-\sum_{i}^{L}\log_{2}p_{h}(x_{i}|x_{<i})/L$. Therefore, an upper bound on the expected BPD error reflects a guarantee on the average performance of the model at the token level, conditioned on previous tokens within the same document, and is a quantity of interest in language modeling.
Table 1: Our best document-level generalization bounds achieved for the GPT-2 architecture for BPD and Top-k token prediction error, all of which are non-vacuous.
@@ 6.2 Non-Vacuous Bounds for GPT-2 Small
Top-1 Error (%) $\mathbf{96.41}$ 100 96.52 100 99.99
Top-10 Error (%) $\mathbf{77.90}$ 84.37 79.36 100 99.98
Top-100 Error (%) $\mathbf{58.34}$ 67.26 75.95 100 99.80
Bits per Dimension $\mathbf{12.12}$ 13.09 14.59 70.76 15.62
6.3 Extending Our Bounds to Larger Models
We use SubLoRA to obtain generalization bounds for much larger variants of GPT-2 of sizes $354$M (GPT-2 medium), $458$M, $773$M (GPT-2 large), and $849$M parameters. Table 2 shows that our simple compression approach yields non-vacuous bounds for models with nearly a billion parameters. Moreover, we see that the smallest model, where we previously performed experiments and tuned our hyperparameters, actually achieves the worst bound on bits per dimension as we scale the models up. In conclusion, our approach extends naturally to much larger language models and proves that it is possible to achieve tighter bounds as we increase the size of the model.
Table 2: Non-vacuous bounds achieved for GPT-2 architectures with different sizes, ranging from $124$ to $849$ million parameters. We report below the bounds on the bits-per-dimension (BPD), Top-1 Error, Top-10 Error, and Top-100 Error. All of the BPD bounds are non-vacuous and tighter than the GPT-2 small bounds.
Model Size BPD Top-1 Error Top-10 Error Top-100 Error
$124$M (GPT-2 small) $12.12$ $96.41$ $77.90$ $58.34$
$354$M (GPT-2 medium) $11.96$ $95.99$ $78.36$ $58.4$
$458$M $11.95$ $96.69$ $78.03$ $58.49$
$773$M (GPT-2 large) $12.10$ $96.17$ $78.57$ $59.25$
$849$M $12.01$ $96.51$ $78.53$ $58.89$
Note 0 (Limitations).
Note that due to computational constraints, we pre-train the larger GPT-2 variants with SubLoRA only for a limited number of hyperparameter settings in contrast to the $124$M model for which we did a thorough hyperparameter sweep. It is likely that the tightest empirically achievable bounds are much stronger for the new large models than what we report in Table 2.
@@ 7 Understanding the Generalization of LLMs
As language models grow in size, it is clear that they gain an increasing capacity to fit their training data. On the one hand, this increasing capacity might mean that, as LLMs become capable of learning increasingly complex functions, they become increasingly likely to merely memorize their training samples and not perform any meaningful generalization beyond their training corpora. After all, they have many more parameters to use in fitting the data. On the other hand, large language models have proven to be surprisingly capable of generalizing, often extending to tasks that seem quite different from the training objective.
We investigate the tension between these two narratives along several fronts: We assess how generalization bounds change with the size of the model, whether language models can form a compression of the training data even when accounting for their large size, and how structure in the training data affects the generalization of the learned model. In Appendix C, we use our bounds to quantify of the benefits of pre-training in LLMs.
@@ 7.1 Larger Models Are More Compressible and Generalize Better
We evaluate our bounds on a collection of LLMs with different numbers of parameters, choosing the appropriate scaling for the width, depth, number of attention heads, etc. Surprisingly, we find that our generalization bounds in fact improve with model size, even as the training dataset is held fixed. With our SubLoRA compression, larger models are more compressible given a fixed training error. These results are shown in Figure 3. While some explanations for why larger models should generalize better have been put forward in the literature (Nakkiran et al., 2021; Gunasekar et al., 2017), the mechanism by which larger models become more compressible is not clear, and we believe this result is noteworthy and requires further investigation.
In addition to constructing generalization bounds, we can use our compressed models to form a compression of the training dataset itself. In Figure 3, we count the number of bits needed to encode the model $C(h)$ and the number of bits to encode the data using the model $C(\{X\}_{i=1}^{m}|h)$, which is the negative log-likelihood of the entire dataset according to the model. Adding these two up, we have a compression of the training dataset using the model, and one which is closely related to our generalization bounds.
@@ 7.2 How Does Generalization of LLMs Depend on Structure in Text?
Neural networks that fit a training dataset of random noise will not be able to generalize, and the ability of overparametrized networks to fit noise implies that uniform convergence is impossible across the general hypothesis class (Nagarajan and Kolter, 2019). This fact is a clear demonstration that the structure of the dataset influences the generalization properties of the model. However, the impact of more subtle structures on generalization is less understood theoretically. Here, we use our bounds to investigate how the temporal order structure relates to generalization.
We train models that explicitly break the temporal structure of the text data by applying random permutations to each sequence during training. Consequently, the model can only make use of the input information as if it were a bag of words. We find that this broken order structure indeed leads to less favorable generalization bounds. Figure 4 shows the best error bounds when the original and perturbed data are used to train the model and evaluate the bounds for the bits per dimension, top-1 error, and top-100 error losses. While the top-1 error bound becomes vacuous as we break the text structure, the top-100 error and bits per dimensions bounds remain non-vacuous. This might be due to the fact that as we perturb the sequence, predicting the next token accurately becomes an extremely difficult task for LLMs, while predicting a token that fits generally into the context, without necessarily being the correct token, is an easier task.
8 Discussion
In this work, we have demonstrated that—despite containing a very large number of parameters—large language models are highly compressible. Using highly compressed LLMs, we were able to compute the first non-vacuous generalization bounds for LLM pretraining. Our findings suggest that the development of tighter compression bounds presents a fruitful avenue for understanding how and why language models generalize. We close with a discussion of the limitations of this work, along with their implications for future generalization theory of language models:
@@ 8 Discussion
Appendix for
Non-Vacuous Generalization Bounds for Large Language Models
Appendix Outline
The appendix is organized as follows:
  • In Appendix A, we provide proofs of the finite hypothesis bound, smoothed prediction bounds for the log-likelihood objective, and the subsampling bounds.
  • In Appendix B, we present our sequence-level bounds for the bits-per-dimension, Top-1 error, Top-10 error and Top-100 error.
  • In Appendix C, we discuss our LLM fine-tuning bounds and highlight the value of pre-training large language models.
  • In Appendix D, we provide the experimental details and hyperparameter selection for our experiments, and describe the document-level empirical risk evaluation.
  • Lastly, in Appendix E we compare the text generated by a GPT-2 small model pre-trained in the standard fashion and a GPT-2 small model pre-trained with SubLoRA.
@@ A.1 Finite Hypothesis Bound
Consider a bounded risk $R(h,x_{i})\in[a,a+\Delta]$ and a finite hypothesis space $h\in\mathcal{H}$ for which we have a prior $P(h)$ that does not depend on $\{x_{i}\}$. Let the empirical risk $\hat{R}(h)=\frac{1}{m}\sum_{i=1}^{m}R(h,x_{i})$ be a sum over independent random variables $R(h,x_{i})$ for a fixed hypothesis $h$. Let $R(h)=\mathbb{E}[\hat{R}(h)]$ be the expected risk.
@@
$R(h)\leq\hat{R}(h)+\Delta\sqrt{\frac{\log 1/P(h)+\log 1/\delta}{2m}},$ (A.1)
@@
$t(h)=\Delta\sqrt{\frac{\log 1/P(h)+\log 1/\delta}{2m}}$ (A.2)
@@
$R(h)\leq\hat{R}(h)+\Delta\sqrt{\frac{\log 1/P(h)+\log 1/\delta}{2m}}.$